OpenStudy (anonymous):

How do i do this? log(3)2c= -2

OpenStudy (anonymous):

3^-2=2c

OpenStudy (anonymous):

solve for c

OpenStudy (anonymous):

how you get 3^-2=2c

OpenStudy (anonymous):

well log is just the inverse function of an exponent logb x = y is b^y=x And plug in your numbers and variables

OpenStudy (anonymous):

\[b=a^{x}\] then \[x=\log_{a}b \] so the power becomes the subject of the equation but we are going back hence \[3^{-2}=2c\]

OpenStudy (anonymous):

so how do i solve for c

OpenStudy (anonymous):

\[c=1/18\]

OpenStudy (anonymous):

how do i solve for c

OpenStudy (anonymous):

divide by 2 both sides and evaluate

OpenStudy (anonymous):

\[2c=1/3^{2}\] \[2c=1/9\] \[c=1/9\div2=1/18\]